OpenStudy (anonymous):

Becca's family has 4 gallons of milk. They drink 5/8 gallon per day. How long will their milk last? A.2 1/2 Days B.4 Days C.4 5/8 D.6 2/5

OpenStudy (tkhunny):

Have you considered division?

OpenStudy (anonymous):

5/ 8 is .625 in decimal form

OpenStudy (tkhunny):

\(\dfrac{4}{5/8} = 4\cdot\dfrac{8}{5} = \dfrac{4}{1}\cdot\dfrac{8}{5} = \dfrac{32}{5}\)

OpenStudy (anonymous):

4 / 0.625 =?

OpenStudy (tkhunny):

Right. Try not to be afraid of fractions, particularly when the answers are given in fractions.

OpenStudy (anonymous):

Which ever way you do it you will get the same answer. My way and @tkhunny will give you the same answer :)

OpenStudy (anonymous):

*@/tkhunny way will

OpenStudy (anonymous):

@arilove1d and @tkhunny thanks that helped a lot :)

OpenStudy (anonymous):

No problem fellow directioner :)